# i-PI 3.0: a flexible, efficient framework for advanced atomistic simulations

### v1

@@ IV.1 Bosonic and Fermionic Path Integral Molecular Dynamics
The new version of i-PI  supports PIMD simulations that include exchange effects, using an algorithm introduced by Hirshberg et al.[89, 90] In this approach, PIMD simulations of bosons scale quadratically,[90] rather than exponentially, with $N$. This is achieved by replacing the standard ring polymer potential with $V_{B}^{(N)}$, which is defined by the recurrence relation,
@@
Here, $E_{N}^{(k)}$ is the spring potential of a ring polymer connecting the particles $N-k+1,\ldots,N$ in a cycle, and $V_{B}^{(0)}=0$. This potential, and the forces it induces, are computed in quadratic time with $N$ and in linear time with the number of beads.[90]
The i-PI implementation allows the efficient simulation of systems composed of $\mathord{\sim}1000$ particles. A comparison of the scalability of bosonic and standard PIMD simulations in i-PI is presented in Fig. 5, using the same normal-mode integrator (bab, see below) in both cases . The bosonic simulation has quadratic scaling while the distinguishable one scales linearly. For $N=64$ particles, the cost of bosonic exchange effects is approximately $12$ times that of standard PIMD, and for $N=1024$ it is only $44$ times slower with the same integrator.
Figure 5: Scaling with the number of particles of the propagation of the spring forces in Cartesian coordinates, for distinguishable and bosonic systems. Both simulations were run with the bab progpagator with nmts=10.
Using bosonic PIMD in i-PI is straightforward, and involves adding the following section to the xml input file:
<normal_modes propagator=’bab’>
<bosons>...</bosons>
<nmts>M</nmts>
</normal_modes>
The bosonic atoms are listed using the bosons tag. This can be done as a list of atoms (zero-based), such as <bosons id=’index’> [0, 1, 2] </bosons>, or by a list of labels, such as <bosons id=’label’> [’He’] </bosons>. At the moment, the integration should be done in Cartesian coordinates (normal_modes propagator=’bab’), because an analytic expression for the normal-modes propagator of the free ring polymer part of the Hamiltonian is known for bosons. Because the frequencies of the springs are typically higher than the physical frequencies, the use of multiple-time stepping in the Cartesian integrator is recommended (via the nmts tag, where M is the factor by which to reduce the timestep).
In Fig. 6, we present the average energy of 512 trapped, non-interacting cold bosons. By comparing with distinguishable particle simulations, the relevance of exchange effects becomes evident.
Figure 6: The energy per particle of $N=512$ trapped non-interacting bosons as a function of temperature (indicated by the dimensionless number $\beta\hbar\omega_{0}$, where $\omega_{0}$ is the harmonic trap constant), with $P=42$ beads. The shaded areas indicate an error region of ±$\pm 0.5\%$.
A way to assess the extent of exchange effects in bosonic simulations is to analyse the following system properties provided within i-PI: (1) exchange_distinct_prob, which is the probability of the ring-polymer configuration corresponding to distinguishable particles:
@@
where $E_{i}^{(1)}$ is the spring potential of the ring polymer connecting the beads of particle $i$ in a separate ring. This quantity tends to $1$ in the regimes where exchange is negligible. (2) exchange_all_prob, which is the probability of the ring-polymer configuration connecting all the bosons, scaled by $1/N$:
@@
Several estimators require modification when performing bosonic PIMD simulations. For example, the primitive estimator for the kinetic energy (kinetic_td) is modified according to Eq. 1 (as shown in Eq. (2) in the SI of  Hirshberg et al. [89]) while the standard virial kinetic energy estimator can be used for bosons in confined systems. On the other hand, the centroid-virial kinetic estimator is not applicable due to the lack of an appropriate centroid definition.
Lastly, fermionic statistics can be obtained by first performing bosonic simulations and then reweighting the results.[91, 111, 112] This is done by using the fermionic sign estimator (the fermionic_sign property of i-PI). For every position-dependent operator, the corresponding fermionic estimator is obtained from the bosonic estimator by multiplying with the fermionic sign, and dividing by the average fermionic sign in the simulation (see Eq. (9) of Ref. [91]). The average sign is always positive but for simulations at low temperatures or with large $N$, where the sign problem is severe, it gets close to zero. Fig. 8 illustrates the behaviour of the fermionic energy estimator for $N=3$ non-interacting, trapped fermions at several temperatures. While the distribution of the bosonic energy estimator throughout the simulation is always positive, the fermionic energy estimator changes sign, its distribution is bi-modal around $E=0$, and more symmetric at lower temperatures. The massive cancellations of positive and negative contributions make the mean fermionic energy harder to converge in comparison to the bosonic case.
Figure 8: The distribution of bosonic and fermionic energy estimator, $P(E)$, for $N=3$ trapped non-interacting particles at two representative temperatures indicated by the dimensionless number $\beta\hbar\omega_{0}$, where $\omega_{0}$ is the harmonic trap constant.
@@ IV.2 Nuclear propagation by additional dynamical variables
i-PI 3.0 can employ numerical quantities provided by the driver (beyond the usual forces and stress arrays) to run advanced types of dynamics. This supplementary communication is achieved through an additional string that can be passed in the extras field of i-PI within the same socket communication. When such string is JSON-formatted, an automatic parser makes the respective quantity available in the code and accessible as a dictionary. More importantly, when the data is numeric, it is added as a property of the force object, guaranteeing a seamless integration with other i-PI features, and becoming available to be used in any type of algorithm. In the next section, we present two examples of advanced simulations that employ electronic structure quantities as “additional” variables to modify the atomic forces and therefore alter the nuclear dynamics.
@@ IV.2.1 Non-adiabatic tunnelling rates in metallic systems
The vanishing band gap in metals allows for nuclear movements to induce electronic excitations, even at relatively low velocities. These effects are not captured by the Born-Oppenheimer approximation and therefore are collectively referred to as non-adiabatic effects (NAEs). NAEs in metals are qualitatively different from the ones found in molecules, due to the continuous manifold of available empty electronic states present in the former, and thus require a different set of theoretical approaches. A frequently used technique to simulate these types of NAEs is to decorate the otherwise Born-Oppenheimer dynamics with an electronic frictional kernel and a stochastic force.[113, 114, 115] The electronic friction formalism has been derived from first principles starting from the exact factorization of the wave function, initially at 0 K,[116, 117] and more recently at finite temperatures.[118] Moreover, in the case of a theory of non-interacting electrons, such as DFT, the friction kernel can be computed explicitly from ab initio calculations.[119]
When studying the kinetics of light atom diffusing in metals, NQEs can become as important as NAEs. To address this complex problem some of the authors have recently derived a theoretical framework that can capture both types of effects.[97] The method called ring polymer instanton with explicit friction (RPI-EF) was derived by combining the semiclassical instanton rate theory in its path integral discretized form,[120] with the ab initio electronic friction formalism. The framework is general enough to incorporate the spatial and frequency dependence of the friction tensor and, owing to the computational efficiency of the parent methods, [121, 122] it is suitable for first-principles calculations of reaction rates in complex, high-dimensional systems.[98]
The RPI-EF method is algorithmically similar to classical Eyring transition rate theory and requires the search of a first-order saddle point. In this case, however, the saddle point is not the reaction transition state but rather an imaginary time trajectory that makes the Euclidean action stationary. This special trajectory is known as instanton and can be interpreted as the main tunnelling pathway.[123] In the limiting case of a position-independent friction kernel (see a more general expression in Ref. [97]) the discretized Euclidean action, $S_{P}$ can be expressed as
@@
where $P$ is the number of discretization points or beads, $N$ is the number of atoms, $\beta$ is the inverse temperature, $\bm{Q}^{(l)}$ represents the free ring-polymer normal mode coordinates, $\omega_{l}$ are the free RP normal mode frequencies, and $\tilde{\eta}$ is the Laplace transform of the (electronic) friction kernel. The RPI-EF calculation consists of a series of geometry optimizations using progressively more discretization points until the converged instanton pathway is achieved. A detailed description of the individual steps involved in RPI and RPI-EF simulations, as well as practical guidelines, can be found elsewhere.[124, 125, 98]
The i-PI  implementation of RPI-EF builds upon the RPI implementation from the previous i-PI  release. The primary distinction regards the additional terms that contain the friction kernel. In the case of a position-dependent friction, these terms require updates at each step of the geometry optimization, which has been made possible by the new JSON string communication described earlier. The friction terms are enabled by setting the friction flag inside the instanton block, frictionSD. Since the theory assumes that the frequency dependence of the friction kernel is position-independent, the user needs to provide the normalized spectral density, $\tilde{\eta}(\omega)$, that can be specified with the flag fric_spec_dens and the energy at which it has been normalized to one by using the flag fric_spec_dens_ener. A minimal example of the section motion to run RPI-EF simulations with i-PI  is
@@
<fric_spec_dens mode=’file >
@@
In Fig. 9, we present the instanton geometry obtained for a double well potential without friction, with constant friction, and with a position-dependent friction. The inclusion of constant friction causes the particle to behave more classically, resulting in the shrinkage of the instanton geometry when compared to the calculation without friction. The position-dependent friction result shows that the instanton geometry avoids regions with large friction, indicating that non-adiabatic effects can steer the main tunnelling pathway and, consequently, steer quantum dynamics.[98]
Figure 9: Instanton pathways obtained in double-well (DW) potential with a) no friction, b) constant friction and c) position-dependent friction. The instanton pathways, potential energy surface and friction profiles, $\tilde{\eta}$, are depicted as red dotted lines, black solid lines and blue dashed lines, respectively. The analytical expression of the potential and friction profiles can be found in Ref. [98].
@@ IV.2.2 Dynamics with time-dependent external field
Many of these phenomena rely on the existence of anharmonic coupling between different vibrational modes that lead to the so-called nonlinear effects in vibrational dynamics.[130, 131] A popular approach to simulate these experiments and capture the associated nonlinear dynamics is based on the electric dipole approximation (EDA). Within EDA, one can further assume that the changes induced by the electric field $\bm{\mathcal{E}}$ on the electronic structure are not significant, and expand the electric enthalpy function of a system around the vanishing-field condition up to first order. This leads to the following expression for the (external field dependent) potential energy
@@
The i-PI implementation for driven vibrational motion is based on the definition of a time-dependent integrator for the equations of motion, called EDAIntegrator, built on top of the already implemented integrators. The role of this integrator is simply to add $\bm{F}^{\text{ext}}(t)$ to the momenta within the Verlet algorithm. This was implemented in a new motion class of i-PI, called DrivenDynamics, which inherits all the attributes and methods of the already implemented Dynamics class.
@@
<motion mode=‘driven_dynamics’>
<driven_dynamics mode=‘eda-nve’>
<timestep units=‘femtosecond‘> 0.3 </timestep>
@@
<amp units=‘atomic_unit’> [0,5e-2,0] </amp>
<freq units=‘thz’> 115 </freq>
<peak units=‘picosecond’> 0.2 </peak>
<sigma units=‘picosecond’> 0.05 </sigma>
@@
As a benchmark example, a single water monomer has been driven using fixed BECs and the Partridge-Schwenke water model.[135] The parameters adopted for the electric field are the ones shown in the input example above. The resulting trajectory, initialized with vanishing velocities has then been projected along the 3 vibrational modes of the system, which have been computed using the harmonic vibrational analysis provided by i-PI. The vibrational modes are shown in Fig. 10a. The oscillating frequency of the field at around 3836 cm-1 (115 THz) was chosen to be resonant with the symmetric-stretch mode. Fig. 10b shows the energy given by the sum of the three harmonic vibrational mode energies ($\rm E_{\rm bend},E_{\rm sym.stretch},E_{\rm asym.stretch}$). $\rm E_{\rm anharm}$ is computed as the difference between the total energy and the sum of the harmonic energies. The profile of the applied field is also shown in Fig. 10. As expected, the resonant vibrational mode is excited and the energy of the system increases while the driving field assumes non-negligible values. Afterwards, the field decays to zero but the system remains excited due to the lack of a dissipation mechanism. By comparing the different components of the energy, it is evident that the harmonic contribution is dominated by the symmetric-stretch mode and that the anharmonicity plays a relatively minor role in this simple example.
Figure 10: (a) Vibrational modes of a water monomer: bending $\rm E_{\rm bend}$, symmetric-stretch $\rm E_{\rm str,~{}sym}$, and asymmetric-stretch $E_{\rm str,~{}asym}$. (b) Contributions to the total energy along a driven trajectory. Time series of the harmonic contribution to the energy of different vibrational modes, the total energy $\rm E_{\rm tot}$, and the anharmonic contribution $\rm E_{\rm anharm}$ of a water monomer along a driven trajectory. The $\rm E_{\rm str,~{}sym}$ mode is resonant with the applied electric field $\bm{\mathcal{E}}$ (shown in purple).

### v2

@@ IV.1 Bosonic and Fermionic Path Integral Molecular Dynamics
Figure 5: Scaling with the number of particles of the propagation of the spring forces in Cartesian coordinates, for distinguishable and bosonic trapped, non-interacting systems. Both simulations were run with the bab propagator with nmts=10. Simulations were run on a node with two AMD EPYC 7763 Processors with 64 cores each, blocking the entire node for every run and using a single i-Pi force-field client.
Version 3.0 of i-PI  supports PIMD simulations that include exchange effects, using an algorithm introduced by Hirshberg et al.[108, 109] In this approach, PIMD simulations of bosons scale quadratically,[109] rather than exponentially, with $N$. This is achieved by replacing the standard ring polymer potential with $V_{B}^{(N)}$, which is defined by the recurrence relation,
@@
Here, $E_{N}^{(k)}$ is the spring potential of a ring polymer connecting the particles $N-k+1,\ldots,N$ in a cycle, and $V_{B}^{(0)}=0$. This potential and the forces it induces are computed in quadratic time with $N$ and in linear time with the number of beads.[109]
The i-PI implementation allows the efficient simulation of systems composed of $\mathord{\sim}1000$ particles. A comparison of the scalability of bosonic and standard PIMD simulations in i-PI is presented in Fig. 5, using the same integrator (bab, see below) in both cases, for a system of trapped non-interacting particles. The bosonic simulation has quadratic scaling while the distinguishable one scales linearly. For $N=64$ particles, the cost of bosonic exchange effects is approximately $12$ times that of standard PIMD, and for $N=1024$ it is only $44$ times slower with the same integrator.
Figure 6: The energy per particle of $N=512$ trapped non-interacting bosons as a function of temperature (indicated by the dimensionless number $\beta\hbar\omega_{0}$, where $\omega_{0}$ is the harmonic trap constant), with $P=42$ beads. The dashed lines indicate the analytical results and the shaded areas indicate an error region of ±$\pm 0.5\%$.
Performing bosonic PIMD in i-PI is straightforward and involves adding the following section to the xml input file: <normal_modes propagator=’bab’> <bosons>...</bosons> <nmts>M</nmts> </normal_modes> The bosonic atoms are listed using the bosons tag. This can be done as a list of atoms (zero-based), such as <bosons id=’index’> [0, 1, 2] </bosons>, or by a list of labels, such as <bosons id=’label’> [’He’] </bosons>. At the moment, the integration should be done in Cartesian coordinates (normal_modes propagator=’bab’), because an analytic expression for the normal-modes propagator of the free ring polymer part of the Hamiltonian is not known for bosons. Because the frequencies of the springs are typically higher than the physical frequencies, the use of multiple-time stepping in the Cartesian integrator is recommended. [133] This is done by setting the nmts tag (standing for the number of iterations for each multiple-time stepping level) to M, where M is the ratio between the time step used for the physical forces and that used for the spring term.[19] In Fig. 6, we present the average energy of 512 trapped, non-interacting cold bosons. By comparing with distinguishable particle simulations, the relevance of exchange effects becomes evident.
A way to assess the extent of exchange effects in bosonic simulations is to analyse the following probabilities, which can be accessed through the properties tag in i-PI: (i) exchange_distinct_prob, which is the probability of the ring-polymer configuration corresponding to distinguishable particles:
@@
where $E_{i}^{(1)}$ is the spring potential of the ring polymer connecting the beads of particle $i$ in a separate ring. This quantity tends to $1$ in the regimes where the exchange is negligible. (ii) exchange_all_prob, which is the probability of the ring-polymer configuration connecting all the bosons, scaled by $1/N$:
@@
Figure 8: The distribution of bosonic and fermionic energy estimator, $P(E)$, for $N=3$ trapped non-interacting particles at two representative temperatures indicated by the dimensionless number $\beta\hbar\omega_{0}$, where $\omega_{0}$ is the harmonic trap constant: (a) high temperature exhibiting mild cancellations between the two peaks of the bimodal distribution of the fermionic energy estimator, and (b) low temperature where the cancellation is near-complete and the sign problem is more severe.
Several estimators require modifications when performing bosonic PIMD simulations. For example, the primitive estimator for the kinetic energy (kinetic_td) is modified according to Eq. 1 (as shown in Eq. 2 in the SI of  Hirshberg et al. [108]) while the standard virial kinetic energy estimator can be used for bosons in confined systems. On the other hand, the centroid-virial kinetic estimator is not applicable due to the lack of an appropriate centroid definition.
Lastly, fermionic statistics can be obtained by first performing bosonic simulations and then reweighting the results.[110, 134, 135] This is done by using the fermionic sign estimator (the fermionic_sign property of i-PI). For every position-dependent operator, the corresponding fermionic estimator is obtained from the bosonic estimator by multiplying with the fermionic sign, and dividing by the average fermionic sign in the simulation (see Eq. (9) of Ref. [110]). The average sign is always positive but for simulations at low temperatures or with large $N$, where the sign problem is severe, it gets close to zero. Fig. 8 illustrates the behaviour of the fermionic energy estimator for $N=3$ non-interacting, trapped fermions at several temperatures. While the distribution of the bosonic energy estimator throughout the simulation is always positive, the fermionic energy estimator changes sign, its distribution is bi-modal around $E=0$, and more symmetric at lower temperatures. The near-complete cancellations of positive and negative contributions make the mean fermionic energy harder to converge in comparison to the bosonic case.
@@ IV.2 Nuclear propagation by additional dynamical variables
i-PI 3.0 can employ numerical quantities provided by the driver (beyond the usual forces and stress arrays) to run advanced types of dynamics. This supplementary communication is achieved through an additional string that can be passed in the extras field of i-PI within the same socket communication. When such string is JSON-formatted it is automatically parsed, and the additional quantities returned by the driver are made available in the code and accessible as a dictionary. More importantly, when the data is numeric, it is added as a property of the force object, guaranteeing a seamless integration with other i-PI features, and becoming available to be used in any type of algorithm. In the next section, we present two examples of advanced simulations that employ electronic structure quantities as “additional” variables to modify the atomic forces and therefore alter the nuclear dynamics.
@@ IV.2.1 Non-adiabatic tunnelling rates in metallic systems
The vanishing band gap in metals allows nuclear movements to induce electronic excitations, even at relatively low velocities.[136, 137] These effects are not captured by the Born-Oppenheimer approximation and, therefore, are collectively referred to as non-adiabatic effects (NAEs). NAEs in metals are qualitatively different from the ones found in molecules due to the continuous manifold of available empty electronic states present in the former, and thus, they require a different set of theoretical approaches. A frequently used technique to simulate these types of NAEs is to decorate the otherwise Born-Oppenheimer dynamics with an electronic frictional kernel and a stochastic force.[138, 139, 140] The electronic friction formalism has been derived from first principles starting from the exact factorization of the wave function, initially at 0 K,[141, 142] and more recently at finite temperatures.[143] Moreover, in the case of a theory of non-interacting electrons, such as DFT, the friction kernel can be computed explicitly from ab initio calculations.[144]
When studying the kinetics of light atom diffusing in metals, nuclear quantum effects (NQEs) can become as important as NAEs. To address this complex problem some of the authors have recently derived a theoretical framework that can capture both types of effects.[113] The method called ring polymer instanton with explicit friction (RPI-EF) was derived by combining the semiclassical instanton rate theory in its path integral discretized form,[145] with the ab initio electronic friction formalism. The framework is general enough to incorporate the spatial and frequency dependence of the friction tensor and, owing to the computational efficiency of the parent methods, [146, 147] it is suitable for first-principles calculations of reaction rates in complex, high-dimensional systems.[114]
The RPI-EF method is algorithmically similar to classical Eyring transition rate theory and requires the search of a first-order saddle point. In this case, however, the saddle point is not the reaction transition state but rather an imaginary time trajectory that makes the Euclidean action stationary. This special trajectory is known as instanton and can be interpreted as the main tunnelling pathway.[148] In the limiting case of a position-independent friction kernel (see a more general expression including the position dependence of the friction kernel in Ref. [113]) the discretized Euclidean action, $S_{P}$ can be expressed as
@@
where $P$ is the number of discretization points or beads, $N$ is the number of atoms, $\beta$ is the inverse temperature, $\bm{Q}^{(l)}$ represents the free ring-polymer normal mode coordinates, $\omega_{l}$ are the free ring-polymer (RP) normal mode frequencies, and $\tilde{\eta}$ is the Laplace transform of the (electronic) friction kernel. $S_{P}$ is related to the potential energy of the ring polymer $U_{P}$ by $S_{P}/\hbar=\beta_{P}U_{P}$. The RPI-EF calculation consists of a series of geometry optimizations on $U_{P}$ using progressively more discretization points until the converged instanton pathway is achieved. A detailed description of the individual steps involved in RPI and RPI-EF simulations, as well as practical guidelines, can be found elsewhere.[149, 150, 114]
Figure 9: Instanton pathways obtained in double-well potential with a) no friction, b) constant friction and c) position-dependent friction. The instanton pathways, potential energy surface and friction profiles, $\tilde{\eta}$, are depicted as red dotted lines, black solid lines and blue dashed lines, respectively. The analytical expression of the potential and friction profiles can be found in Ref. [114].
The i-PI  implementation of RPI-EF builds upon the RPI implementation from the previous i-PI  release. The primary distinction regards the additional terms that contain the friction kernel. In the case of position-dependent friction, these terms require updates at each step of the geometry optimization, which has been made possible by the automatic parsing of JSON-formatted data returned by the client. The friction terms are enabled by setting the friction flag inside the instanton block, frictionSD. Since the theory assumes that the frequency dependence of the friction kernel is the same for all positions, the user needs to provide the normalized spectral density, $\tilde{\eta}(\omega)$, that can be specified with the flag fric_spec_dens and the energy at which it has been normalized to one by using the flag fric_spec_dens_ener. A minimal example of the section motion to run RPI-EF simulations with i-PI  is
@@
<fric_spec_dens mode=’file’>
@@
In Fig. 9, we present the instanton geometry obtained for a 1D double well potential without including any friction, including a constant friction profile, and including an analytical position-dependent friction profile. The inclusion of constant friction causes the particle to behave more classically, resulting in the shrinkage of the instanton geometry when compared to the calculation without friction. The result with position-dependent friction shows that the instanton geometry avoids regions with large friction. As discussed in Refs. [114, 151], this indicates that non-adiabatic effects can steer the main tunnelling pathway and, consequently, steer quantum dynamics.
@@ IV.2.2 Dynamics with time-dependent external field
Figure 10: (a) Vibrational modes of a water monomer: bending $\rm E_{\rm bend}$, symmetric-stretch $\rm E_{\rm str,~{}sym}$, and asymmetric-stretch $E_{\rm str,~{}asym}$. (b) Contributions to the total energy along a driven trajectory. Time series of the harmonic contribution to the energy of different vibrational modes, the total energy $\rm E_{\rm tot}$, and the anharmonic contribution $\rm E_{\rm anharm}$ of a water monomer along a driven trajectory. The $\rm E_{\rm str,~{}sym}$ mode (blue) is resonant with the applied electric field $\bm{\mathcal{E}}$ (shown in purple). The contributions from $\rm E_{\rm str,~{}asym}$ (yellow) and $\rm E_{\rm bend}$ (red) remain close to zero throughout.
Many of these phenomena rely on the existence of anharmonic coupling between different vibrational modes that lead to the so-called nonlinear effects in vibrational dynamics.[156, 157] A popular approach to simulate these experiments and capture the associated nonlinear dynamics is based on the electric dipole approximation (EDA). Within EDA, one can assume that the changes induced by the electric field $\bm{\mathcal{E}}$ on the electronic structure are not significant, and expand the electric enthalpy function of a system around the vanishing-field condition up to first order. This leads to the following expression for the (external field dependent) potential energy
@@
The i-PI implementation for driven vibrational motion is based on the definition of a time-dependent integrator for the equations of motion, called EDAIntegrator, built on top of the existing integrators. The role of this integrator is simply to add $\bm{F}^{\text{ext}}(t)$ to the momenta within the Verlet algorithm. This was implemented in a new motion class of i-PI, called DrivenDynamics, which inherits all the attributes and methods of the Dynamics class.
@@
<motion mode=’driven_dynamics’>
<driven_dynamics mode=’eda-nve’>
<timestep units=’femtosecond‘> 0.3 </timestep>
@@
<amp units=’atomic_unit’> [0,5e-2,0] </amp>
<freq units=’thz’> 115 </freq>
<peak units=’picosecond’> 0.2 </peak>
<sigma units=’picosecond’> 0.05 </sigma>
@@
As a benchmark example, a single water monomer has been driven using fixed BECs and the Partridge-Schwenke water model.[161] The parameters adopted for the electric field are the ones shown in the input example above. The resulting trajectory, initialized with vanishing velocities has then been projected along the 3 vibrational modes of the system, which have been computed using the harmonic vibrational analysis provided by i-PI. The vibrational modes are shown in Fig. 10a. The oscillating frequency of the field at around 3836 cm-1 (115 THz) was chosen to be resonant with the symmetric-stretch mode.
Fig. 10b shows the energy given by the sum of the three harmonic vibrational mode energies ($\rm E_{\rm bend},E_{\rm sym.stretch},E_{\rm asym.stretch}$). $\rm E_{\rm anharm}$ is computed as the difference between the total energy and the sum of the harmonic energies. The profile of the applied field is also shown in Fig. 10. As expected, the resonant vibrational mode is excited and the energy of the system increases while the driving field assumes non-negligible values. Afterwards, the field decays to zero but the system remains excited due to the lack of a dissipation mechanism. By comparing the different components of the energy, it is evident that the harmonic contribution is dominated by the symmetric-stretch mode and that in this simple example the anharmonicity plays a relatively minor role.